3a-15+19=-2
can you please help me

Answer by ankor@dixie-net.com(22740)

3a - 15 + 19 = -2
do the math
3a + 4 = -2
subtract 4 from both sides
3a + 4 - 4 = -2 - 4
do the math
3a = -6
Divide both sides by 3
a = -6/3
a = -2
:
:
Check this in the original problem, replace a with -2
3(-2) - 15 + 19 = -2
-6 - 15 + 19 = -2
-2 = -2, confirms our solution of a = -2